. If the lever arm of a screw is 18 inches long and the pitch of the threads is 0.125


inch, how much weight will the jack lift if 50 pounds of force is applied at the end


of the lever arm ignoring friction?

1
Expert's answer
2021-11-14T17:15:05-0500

The weight is


"W=F·\frac{2\pi R}{p}=50·\frac{2\pi 18}{0.125}=45.2·10^3\text{ lb}."